if f(x)=1-2x, find f(f(2))
Answer by Theo(13342)

you are given that f(x) = 1 - 2x.

f(2) would be found by replacing x with 2.

you would get f(2) = 1 - 2(2) = 1-4 = -3

f(f(2)) would be found by replacing x with f(2) which is equal to -3.

you would get f(f(2)) = f(-3) = 1 - 2(-3) = 1+6 = 7.

you work your way from the inner set of parentheses out.

you first find f(2) and then you find f(f(2)).

this is no different than a composite function of gof(x), where f(x) = 1 - 2x and g(x) = 1 - 2x.

you would first find f(2) = 1 - 2(2) = -3

you would then find g(f(2)) = g(-3) = 1 - 2(-3) = 1+6 = 7

here'a a reference on composite functions.

http://www.purplemath.com/modules/fcncomp2.htm